OpenStudy (anonymous):

Find the equation of the tangent line to the curve y=√x at x=4. please explain step by step? thank you!! :)

OpenStudy (dumbcow):

the derivative at the point (4,2) is the slope of tangent line. then use point-slope form to get equation of line y-y1 =m(x-x1)

OpenStudy (anonymous):

so y-2 = m(x-4) ? and how would we find m ?

OpenStudy (dumbcow):

slope = derivative

OpenStudy (anonymous):

okay :) would we use f(a+h)-f(a) --------- h ??

OpenStudy (dumbcow):

you could...have you learned the power rule for finding derivatives of polynomials ?

OpenStudy (anonymous):

yes :) f(x)=x^n f ' (x) = nx^(n-1) right? :)

OpenStudy (dumbcow):

\[\frac{d}{dx} x^n = n x^{n-1}\]

OpenStudy (dumbcow):

haha yeah use that

OpenStudy (agent0smith):

\[\Large y = √x = x^{1/2}\]

OpenStudy (anonymous):

haha okie :P so we would have y=√x = x^(1/2) = (1/2) * x^((1/2)-(-.5) ?

OpenStudy (agent0smith):

^idk what you're doing with the exponent. Reduce it by 1 (subtract 1 from it)

OpenStudy (anonymous):

oh oops :p okay so y=√x = x^(1/2) = (1/2) * x^((1/2)-1) = (1/2) * x^-1/2 ? =(1/2)x ------- x^(1/2) ??

OpenStudy (agent0smith):

Why's there an extra x...

OpenStudy (anonymous):

oh oops typo haha so = x/x^(1/2) ?

OpenStudy (agent0smith):

Still an extra x. This is right: = (1/2) * x^-1/2 so\[\Large y ' = \frac{ 1 }{2 \sqrt x }\]

OpenStudy (anonymous):

ohh okay... so that's the final equation? is there any way you can write out what happens from the first step to the last step? i'm kinda confused on where they connect :(

OpenStudy (dumbcow):

Given f(x) and point (x1,y1) equation of tangent line is: \[y - y_1 = f'(x_1) (x - x_1)\]

OpenStudy (anonymous):

ohh okay.. so then the points are (4,2) and then we get y=√x = x^(1/2) = (1/2) * x^((1/2)-1) = (1/2) * x^-1/2 = y ' = 1/ (2√x) ? and that is the final equation of the tangent line to the curve y=√x at x=4? did i miss anything? :/

OpenStudy (agent0smith):

That is not the tangent line. That y' gives you the slope, m, in your tangent line: y = mx + b

OpenStudy (agent0smith):

Plug in x=4 to find the slope.

OpenStudy (anonymous):

ohh okay but those steps above are correct so far?

OpenStudy (anonymous):

okay, so 2=(1/(2√x)) * (4) + b ?

OpenStudy (agent0smith):

Find m FIRST using y ' = 1/ (2√x) and x=4

OpenStudy (anonymous):

ohh i see.. so y' = 1/ 4 ?

OpenStudy (anonymous):

oh and is all this using the definition of the slope of the tangent line? i'm supposed to do learn it via definition right now, not shortcut :/ just wondering hahaa

OpenStudy (agent0smith):

Idk if this is really a shortcut. Now you have m, in y = mx+b and you have the point (4,2). Find b.

OpenStudy (anonymous):

haha okie :P so 2=(1/4)(4) + b = 2=1 +b = b=1 ?

OpenStudy (agent0smith):

Yep, now write the line y=mx+b

OpenStudy (anonymous):

\[y=\frac{ 1 }{ 4 }x + 1 \] ?? is that right?

OpenStudy (agent0smith):

Yep

OpenStudy (anonymous):

awesome! so that's the final equation of the tangent line? ?:O

OpenStudy (agent0smith):

Yep. It's the same process every time. Find the slope at the point (using the derivative). Find the line y=mx+b using that slope, and the point it passes through.

OpenStudy (anonymous):

ahh okay awesome!! and also wondering, would you be able to solve this with the f(a+h) - f(a) ---------- h equation??

OpenStudy (anonymous):

or is this the only way to solve for the equation?

OpenStudy (agent0smith):

That just is the long way of finding the slope, m.

OpenStudy (anonymous):

ohh okay... just for future reference, how would i find the slope the long way?

OpenStudy (anonymous):

would i be plugging in like this? \[\lim_{h \rightarrow 0} \frac{ (f(4+h)-f(4) }{ h }\] ??? ;/ @agent0smith

OpenStudy (agent0smith):

Yeah, I think that's right.

OpenStudy (anonymous):

okay:) what would happen next? ;/

OpenStudy (agent0smith):

You'd simplify it until you have the slope, 1/4

OpenStudy (anonymous):

okay cool, thank you!!!